# Zens, Díaz, Durante and Patacchini

### v1

@@ Abstract
In this article, we proposed a low-rank bilinear autoregressive model to uncover interpretable dependence structure within and across the three dimensions (time, space, crime categories) of observed criminal activity tensors, with a specific focus on report data from the city Chicago. Besides disentangling and facilitating inference on complex interdependence patterns over several dimensions, the proposed formulation crucially achieves also a predictive accuracy that is competitive with black-box strategies, thereby providing a realistic, yet interpretable, model-based representation of crime phenomena over space, time and different categories. The results in the application confirm these advantages, and further stress the importance of developing joint models that maintain a systemic perspective on crime phenomena by formally accounting for dependencies across and within all the tree dimensions of a generic criminal activity tensor. This perspective not only contributes to enlarging available knowledge on the mechanisms and determinants of crime, but also facilitates the design and evaluation of improved policies and intervention strategies.
@@
The above advancements motivate also several directions for future research. First, extending the model to incorporate nonlinearities could provide further insights into the complex dynamics of criminal activities that might also yield further gains in predictive accuracy. Relatedly, another interesting direction for future research would be to explicitly incorporate additional structural information, such as knowledge on activity patterns of organized criminal groups or spatial structure, into the modeling framework. The latter could be achieved by introducing external covariates or by incorporating latent spatio-temporal dependencies into the model parameters, leveraging, e.g., results in Hsu et al. (2021). Finally, given the typical dominance of the main diagonal in the estimated matrices ${\bf A}$ and ${\bf B}$, future research could consider other alternatives to the factorization in (2), such as, for example ${\bf A}={\bf D}_{A}+{\bf L}_{A}{\bf Z}_{A}$ and ${\bf B}={\bf D}_{B}+{\bf L}_{B}{\bf Z}_{B}$, respectively, with ${\bf D}_{A}$ and ${\bf D}_{B}$ diagonal. In terms of applications, it would be interesting to develop quantitative strategies for designing and assessing intervention policies guided by the output of the proposed model, especially with reference to the magnitude and structure of potential spillover effects. Expanding the analysis to other crime datasets from different cities and regions would also be important to obtain a more comprehensive perspective on criminal activities in urban environments beyond the city of Chicago. Moreover, the proposed model and analyses can also be applied beyond crime, whenever data are available in the form of activities/choices/behaviors/events over space and time with respect to a number of categories. An example could be employment data in urban environments over time and districts, with respect to different job categories.
@@ 7 Acknowledgments
Carlos Díaz gratefully acknowledges financial support from the National Agency of Research and Development’s (ANID) Millennium Science Initiative Program (NCS2024 058). Daniele Durante is funded by the European Union (ERC, NEMESIS, project number: 101116718). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them.

### v2

@@ Abstract
In this article, we proposed a low-rank bilinear autoregressive model to uncover interpretable dependence structure within and across the three dimensions (time, space, crime categories) of observed criminal activity tensors, with a specific focus on report data from the Chicago city. Besides disentangling and facilitating inference on complex interdependence patterns over several dimensions, the proposed formulation crucially achieves also a predictive accuracy that is competitive with black-box strategies, thereby providing a realistic, yet interpretable, model-based representation of crime phenomena over space, time and different categories. The results in the application confirm these advantages, and further stress the importance of developing joint models that maintain a systemic perspective on crime phenomena by formally accounting for dependencies across and within all the three dimensions of a generic criminal activity tensor. This perspective not only contributes to enlarging available knowledge on the mechanisms and determinants of crime, but also facilitates the design and evaluation of improved policies and intervention strategies.
@@
The above advancements motivate several directions for future research. First, extending the model to incorporate nonlinearities could provide further insights into the complex dynamics of criminal activities that might also yield further gains in predictive accuracy. Relatedly, another interesting direction would be to explicitly incorporate additional structural information, such as knowledge on activity patterns of organized criminal groups or spatial structure, into the modeling framework. The latter could be achieved by introducing external covariates or by incorporating spatio-temporal dependencies into the model parameters, leveraging, e.g., results in Martinez-Beneito et al. (2017) and Hsu et al. (2021), among others. Although the proposed model learns both distance-related effects and departures from pure proximity patterns, additional empirical comparisons against these spatio-temporal alternatives could further illustrate the potential of our formulation. Finally, given the typical dominance of the main diagonal in the estimated matrices $\bf A$} and $\bf B$}, future research could consider alternatives to the factorization employed in (2), such as, for example ${\bf A}={\bf D}_{A}+{\bf L}_{A}{\bf Z}_{A}$ and ${\bf B}={\bf D}_{B}+{\bf L}_{B}{\bf Z}_{B}$, respectively, with ${\bf D}_{A}$ and ${\bf D}_{B}$ diagonal. In terms of applications, it would be interesting to develop quantitative strategies for designing and assessing intervention policies guided by the output of the proposed model, especially with reference to the magnitude and structure of potential spillover effects. Expanding the analysis to other crime datasets from different cities and regions would also be important to obtain a more comprehensive perspective on criminal activities in urban environments beyond Chicago.
@@ 7 Acknowledgments
Carlos Díaz gratefully acknowledges financial support from the National Agency of Research and Development’s (ANID) Millennium Science Initiative Program (NCS2024_058). Daniele Durante is funded by the European Union (ERC, NEMESIS, project number: 101116718). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them.
1
Supplementary Material to “Low-Rank Bilinear Autoregressive Models for Three-Way Criminal Activity Tensors”
8 S1. Reclassification of crimes in the category Miscellaneous Other
Table 2 clarifies how crimes within the category Miscellaneous Other have been redistributed based on the corresponding Illinois Uniform Crime Reporting (iucr) code.
Table 2: Redistribution of Miscellaneous Other (FBI Code 26) by IUCR
FBI Code Target category IUCR codes moved from 26 Notes
09 Arson 1030, 1035, 5003 Existing FBI classification
16 Prostitution 1050 Existing FBI classification
- Criminal trespassing 1330, 1335, 1350, 1360, 1365 New category
- Kidnapping & Child Harm 1537, 1710, 1715, 1725, 1755, 1780, 1790, 1792, 4210, 4220, 4230, 4240, 4255 New category
18 Drug abuse 2091, 2092, 2093, 2111, 2120, 2160 Existing FBI classification
- Threat & harassment 2820, 2825, 2826, 3960, 3966, 3970, 3975, 3980, 4386, 4387 New category
24 Disorderly conduct 2850, 2851, 2890, 2895, 3000, 3200, 3300, 3400, 3770, 3800, 3910, 3920 Existing FBI classification
05 Burglary 4310, 4860, 5007, 5008 Existing FBI classification
Notes: Recodes apply only where the original FBI Code equals 26 (9.8% of total criminal events).